OpenStudy (anonymous):

Find the common solution. 2x + y = 7 5x + y = 9 (?,?)

OpenStudy (anonymous):

Subtract the first equation from the second: 3x = 2 so x = 2/3 and substitute this in the second equation 10/3 + y = 9 so y = 17/3